# Unchecked and Overlooked: Addressing the Checkbox Blind Spot in Large Language Models with CheckboxQA

### v1

@@ 3.5 Baseline Approaches and Human Performance
Open-source LVLMs are evaluated using vLLM [9]. For Qwen 2 VL [24], we reimplemented vLLM’s Multimodal Rotary Position Embedding (M-RoPE) to include dynamic scaling, enabling context lengths of up to 128k tokens. Default vLLM inference engine options were used.
@@ 4 Results and Analysis
We conducted experiments with state-of-the-art commercial and open-source solutions, including models from GPT-4o [16], Gemini 2.0 [6], Qwen 2/2.5 VL [24], and Pixtral [2] families to evaluate how well large vision-language models handle the fine-grained task of interpreting checkboxes.
@@ 4.1 Quantitative Analysis
Table 1 reports the performance of various LVLMs and a human baseline. Among the tested systems, Qwen 2.5 VL 72B attains the highest score at 83.2%, significantly outperforming GPT-4o, similarly to the previous Qwen VL generation. Smaller Qwen variants, Pixtral 12B, and the Gemini series exhibit more modest results in the range of 43.6% to 71.9%. GPT-4o mini remains at the lower end with a score of 40.4%.

### v2

@@ 3.5 Baseline Approaches and Human Performance
Open-source LVLMs are evaluated using vLLM [9] with default inference options.
@@ 4 Results and Analysis
We conducted experiments with state-of-the-art commercial and open-source solutions, including models from GPT-4o [16], Gemini 2.0 [6], Qwen 2.5 VL [24], and Pixtral [2] families to evaluate how well large vision-language models handle the fine-grained task of interpreting checkboxes.
@@ 4.1 Quantitative Analysis
Table 1 reports the performance of various LVLMs and a human baseline. Among the tested systems, Qwen 2.5 VL 72B attains the highest score at 83.2%, significantly outperforming GPT-4o. Smaller Qwen variants, Pixtral 12B, and the Gemini series exhibit more modest results in the range of 43.6% to 71.9%. GPT-4o mini remains at the lower end with a score of 40.4%.